# Understanding the chemistry of temperate exoplanets atmospheres through experimental and numerical simulations

### v1

@@ Acknowledgements.
Data availability. All data can be downloaded from Zenodo: https://doi.org/10.5281/zenodo.17868179.
@@ Appendix A Experimental identification of species
Figure 11: Mass spectrum of an irradiated gas mixture of 95% H2 + 5% CH4 (Gray). Fragmentation patterns (FP) of the 2-carbon atoms (11(a)) and 3-carbon atoms (11(b)) compounds (Pink for the alkanes, orange for the alkenes, green for the alkynes). Each FP is normalized to the intensity of the experimental spectrum at the m/z value corresponding to the most intense peak of the FP. 95% uncertainty bars are displayed in black.
@@
Table 2: Features of products identified by Mass Spectrometry (MS) and InfraRed Spectroscopy (IR).

### v2

@@ 6 Perspectives
Data availability. Mass spectra, IR spectra, and 0D model data can be downloaded from Zenodo: https://doi.org/10.5281/zenodo.17868179.
@@ Appendix A Experimental identification of species
Figure 11: Mass spectrum of an irradiated gas mixture of 95% H2 + 5% CH4 (Gray). Fragmentation patterns (FP) of the two-carbon atoms (11(a)) and three-carbon atoms (11(b)) compounds (Pink for the alkanes, orange for the alkenes, green for the alkynes). Each FP is normalized to the intensity of the experimental spectrum at the m/z value corresponding to the most intense peak of the FP. 95% uncertainty bars are displayed in black.
@@
Table 2: Features of products identified by MS and IR spectroscopy.